OpenStudy (anonymous):

The third angle of an isosceles triangle is 16 degree less than the sum of the other two angles.Find three angles of triangle?

OpenStudy (anonymous):

isoceles triangle has two base angles equal, hence we can consider the base angles to be equal to x degrees each , therefore according to the problem x+x + (2x-16) = 180 ( sum of three angles of a triangle is 180 degrees) or 4x = 180+16 or 4x = 196 or x = 49 degrees . angles are 49 degrees + 49 degrees+ 82 degrees

OpenStudy (anonymous):

Thanku! Very Much Dear.I Got It